# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ I.2 Molecular Captioning
Instruction: What is the HOMO-LUMO gap of this molecule? The compound SELFIES sequence is: ¡SELFIES_compound¿: [N][=C][O][C][=C][C][=Branch1][Ring2][=N][Ring1][=Branch1][C][#N] Property: 0.1487
J.7 Catalyst Prediction
The catalyst prediction task focuses on identifying the appropriate catalysts required to facilitate a given chemical reaction. Catalysts play a crucial role in modifying reaction pathways, lowering activation energy, and improving reaction efficiency without being consumed in the process. The input follows the SELFIES representation, a robust molecular encoding system designed for computational applications. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using “>>” (e.g., “reactant1.reactant2>>product”). This structured representation allows the model to process the reaction as a whole and infer the most suitable catalyst that enables the transformation. One of the primary challenges in catalyst prediction is understanding the nuanced role that catalysts play in different reaction mechanisms. Unlike reagents, which directly participate in the reaction, catalysts provide alternative pathways to enhance reaction kinetics. The prompt template is as follows.
@@
Instruction: Given this chemical reaction, what are some catalysts that could have been used? The reaction is ¡SELFIES_reactants¿: [C][C][=C][C][=C][Branch1][Ring1][C][#N][C][=C][Ring1][Branch2][C][Branch1][C][F][Branch1][C][F][F].[O][=C][C][C][C][=Branch1][C][=O][N][Ring1][=Branch1][Br] ¡SELFIES_product¿: [N][#C][C][=C][C][=C][Branch1][Ring1][C][Br][C][Branch1][=Branch2][C][Branch1][C][F][Branch1][C][F][F][=C][Ring1][N] ¡SELFIES_catalysts¿:[O][=C][Branch1][#C][O][O][C][=Branch1][C][=O][C][=C][C][=C][C][=C][Ring1][=Branch1][C][=C][C][=C][C] [=C][Ring1][=Branch1]
J.8 Solvent Prediction
The solvent prediction task focuses on identifying the appropriate solvents required for a given chemical reaction. Solvents play a crucial role in determining reaction efficiency, influencing factors such as solubility, reaction kinetics, selectivity, and stability of intermediates. To ensure a structured and machine-readable representation, the input follows the SELFIES format, a robust molecular encoding system designed for computational applications. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using the reaction separator “>>” (e.g., “reactant1.reactant2>>product”). This structured format allows the model to interpret the reaction context and infer the most suitable solvents required to facilitate the transformation. One of the key challenges in solvent prediction is understanding the diverse roles solvents play in different reaction mechanisms. The prompt template is as follows.
@@
Instruction: Please propose potential solvents that might have been utilized in the provided chemical reaction. The reaction is ¡SELFIES_reactants¿: [N][#C][C][=C][C][=C][Branch1][C][F][C][=C][Ring1][#Branch1].[O][C][C][C][N][C][Ring1][Branch1] ¡SELFIES_product¿: [N][#C][C][=C][C][=C][Branch1][O][N][C][C][C][Branch1][C][O][C][Ring1][=Branch1][C][=C][Ring1][N] ¡SELFIES_solvents¿: [O]
J.9 Yield Regression
The yield regression task focuses on estimating the proportion of the actual product obtained in a chemical reaction relative to its theoretical maximum. Reaction yield is a critical metric in organic synthesis, pharmaceutical manufacturing, and industrial chemistry, as it directly influences process efficiency, resource utilization, and cost-effectiveness. The input follows the SELFIES format, a robust molecular encoding system tailored for computational chemistry. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using the reaction separator “>>” (e.g., “reactant1.reactant2>>product”). This structured format provides a standardized input for the model, allowing it to interpret the reaction context and estimate the expected yield. One of the key challenges in yield prediction is capturing the complex interplay between reaction conditions, molecular stability, steric effects, and solvent or catalyst influences. The prompt template is as follows.
@@
Instruction: Please propose potential solvents that might have been utilized in the provided chemical reaction. The reaction is ¡SELFIES_reactants¿: [F][C][Branch1][C][F][Branch1][C][F][C][=C][C][=C][Branch1][C][Cl][C][=C][Ring1][#Branch1].[C][C][=C][C][=C][Branch1] [C][N][C][=C][Ring1][#Branch1].[O][=S][=Branch1][C][=O][Branch2][Ring1][=Branch1][O][Pd][N][C][=C][C][=C][C][=C][Ring1][=Branch1][C][=C][C][=C][C][=C][Ring1][=Branch1][Ring1][=C][C][Branch1][C][F][Branch1][C][F][F].[C][O][C][=C][C][=C][Branch1][Ring1][O][C][C][Branch2][Ring2][=N][P][Branch2][Ring1][Branch1][C][C][C][C][C][Branch1][O][C][C][Branch1][Ring2][C][Ring1][=Branch1][C][Ring1] [=Branch2][C][Ring1][#Branch2][C][C][C][C][C][Branch1][O][C][C][Branch1][Ring2][C][Ring1][=Branch1][C][Ring1][=Branch2][C][Ring1][#Branch2] [=C][Ring2][Ring1][=N][C][=C][Branch1][=Branch1][C][Branch1][C][C][C][C][=C][Branch1][=Branch1][C][Branch1][C][C][C][C][=C][Ring1][N][C][Branch1][C][C][C].[C][N][C][C][C][N][C][C][C][N][=C][Ring1][#Branch2][Ring1][=Branch1].[C][C][O][C][=Branch1][C][=O][C][C] [=C][Branch1][C][C][O][N][=Ring1][=Branch1] ¡SELFIES_product¿: [C][C][=C][C][=C][Branch2][Ring1][Ring2][N][C][=C][C][=C][Branch1][=Branch2][C][Branch1][C][F][Branch1][C][F][F][C] [=C][Ring1][#Branch2][C][=C][Ring1][P] Property: 0.1449
J.10 SCF Energy Prediction
The SCF energy prediction task involves determining the self-consistent field (SCF) energy for a given compound. The input is the SELFIES representation of the compound, and the model is tasked with predicting the molecule’s SCF energy. The objective is to provide a reliable basis for understanding the compound’s total electronic energy, which is critical for assessing its stability, reactivity, and potential applications in computational chemistry. The prompt template is as follows.
  Template
[…]
Instruction: Please provide the SCF Energy value for this molecule. If uncertain, provide an estimate. Respond with the numerical value only. The molecule SELFIES sequence is: ¡SELFIES_compound¿:[C][C][C][C][C][C][=Branch1][C][=O][C][=C][C][C][Branch2][Ring1][Ring2][C][C][=Branch1][C][=O][C][Ring1] [=Branch1][C][C][C][C][C][C][C][=Branch1][C][=O][O][O] Property: The SCF Energy for the input molecule is -1.988e+05 eV.
J.11 LogP Prediction
@@
Instruction: I am interested in the LogP of this molecule, could you tell me what it is? If uncertain, provide an estimate. Respond with the numerical value only. The molecule SELFIES sequence is: ¡SELFIES_compound¿: [C][O][C][=C][C][=Branch1][=Branch2][=C][C][=Branch1][Ring2][=C][Ring1][=Branch1][O][C@H1][C@H1][C@@H1] [Branch1][Branch1][C][O][Ring1][Branch1][C@@H1][Branch1][=Branch1][O][C][Ring1][#Branch1][=O][C][=C][C][=Branch1][=C][=C][Branch1] [=Branch2][C][=Branch1][Ring2][=C][Ring1][=Branch1][O][C][O][C][O] Property: The LogP for the input molecule is 2.00.
J.12 Molecular Weight Prediction
@@
Instruction: Please provide me with the Molecular Weight value of this molecule. Determine the Molecular Weight value of this molecule. If uncertain, provide an estimate. Respond with the numerical value only. The molecule SELFIES sequence is: ¡SELFIES_compound¿: [C][O][C][=C][Branch1][#Branch1][C][=C][C][=N][Ring1][=Branch1][C][=Branch1][C][=O][N][C][C][C][C][C][N][Branch1] [Branch1][C][C][Ring1][=Branch1][S][=Branch1][C][=O][=Branch1][C][=O][N][C][=Branch1][C][=O][N][C][C][C][C][C][C][Ring1] [Branch1][C][=C][Ring1][Branch1] Property: The Molecular Weight for the input molecule is 491.60 g/mol.
J.13 Topological Polar Surface Area Prediction
@@
Instruction: I would like to know the Topological Polar Surface Area of this molecule, can you provide it? If uncertain, provide an estimate. Respond with the numerical value only. The compound SELFIES sequence is: ¡SELFIES_compound¿:[C][C][=Branch2][=Branch1][=Branch2][=C][C][O][C][C][Branch1][O][C][Branch1][Ring2][O][Ring1][=Branch1] [Branch1][C][C][C][O][C][C][C][C][Branch2][Ring2][#C][C][Branch2][Ring2][#Branch2][C][Branch2][Ring1][=Branch1][C][C][Branch1][P][C][Ring1] [=Branch1][Branch1][O][C][Ring1][#Branch2][C][Ring2][Ring1][C][O][Ring1][Ring1][O][O][C][C][=C][Ring1][=N][N][C][=C][C][=C][C][=C][Ring1][=Branch2][Ring1][=Branch1][C][C][C] Property: The Topological Polar Surface Area for the input molecule is 96.50 Ų.
J.14 Complexity Prediction
The complexity prediction task focuses on determining the structural complexity of a given chemical compound, a key property that reflects its architectural intricacy, stereochemical richness, and molecular connectivity. Molecular complexity plays a crucial role in synthetic chemistry, drug discovery, and materials science, where it impacts synthetic feasibility, resource requirements, and overall manufacturability. Compounds with high complexity may require multiple synthetic steps, specialized reagents, and intricate reaction conditions, whereas simpler molecules are generally easier to produce and optimize for industrial applications. The input is the SELFIES representation of the compound, and the model is tasked with predicting the compound’s complexity. The objective is to shed light on the molecule’s structural intricacy, which can influence its synthetic accessibility, resource requirements, and overall feasibility in various chemical processes. The prompt template is as follows.
  Template
[…]
Instruction: I need to know the Complexity of this molecule, could you please provide it? If uncertain, provide an estimate. Respond with the numerical value only. The compound SELFIES sequence is: ¡SELFIES_compound¿: [C][C@@H1][Branch2][Ring2][#Branch2][C][C@@H1][Branch1][P][C][=Branch1][C][=O][N][C@@H1][Branch1][C][C][C] [=N][C][=C][S][Ring1][Branch1][N][Branch1][C][C][C][=Branch1][C][=O][C][C@H1][Branch1][C][C][C][Branch1][C][Cl][Branch1][C][Cl][Cl] [C][Branch1][C][Cl][Branch1][C][Cl][Cl] Property: The Complexity for the input molecule is 572.00.
J.15 Description Q&A
@@
Instruction: What is the main component of Lobaplatin that gives it its anticancer properties? The compound SELFIES sequence is: ¡SELFIES_compound¿: [C][C@@H1][Branch1][=Branch1][C][=Branch1][C][=O][O-1][O-1].[C][C][C@H1][Branch1][=Branch1][C@@H1][Ring1] [Ring2][C][N][C][N].[Pt+2] Description: The main component of Lobaplatin that gives it its anticancer properties is the platinum cation (Pt+2).
J.16 Experimental Procedure Prediction
The experimental procedure prediction task focuses on generating a detailed, step-by-step protocol for conducting chemical experiments based on a given set of experimental components. This task plays a crucial role in automated synthesis planning, high-throughput experimentation, and robotic chemistry, where structured experimental procedures are essential for reproducibility, efficiency, and accuracy. The input is a mapping between experimental components and their corresponding SELFIES representations (e.g., “Reactants: \$index\$: SELFIES” …). The model is tasked with producing a structured sequence of operations that associates each component with the detailed steps of the experiment. The objective is to automate the chemical synthesis process by providing executable, structured experimental procedures. The prompt template is as follows.
@@
Instruction: From the provided starting materials and target compound, generate the step-by-step experimental protocol. The Action Sequence: ¡IDX_Reactants_MAP¿:Reactants:\$4\$:[START_SELFIES][O][=C][C][=C][C][=C][Branch1][O][O][C][C][C][O][C][C][O][Ring1] [Branch1][C][=C][Ring1][=C][END_SELFIES] \$3\$: [START_SELFIES][C][C][Branch1][C][C][Branch1][C][C][O-1].[K+1][END_SELFIES] ¡IDX_Product_MAP¿:Product:\$-1\$:[START_SELFIES][C][=C][C][=C][C][=C][Branch1][O][O][C][C][C][O][C][C][O][Ring1][Branch1][C] [=C][Ring1][=C][END_SELFIES] ¡IDX_Catalysts_MAP¿:Catalysts:\$1\$:[START_SELFIES][C][P+1][Branch1][=Branch2][C][=C][C][=C][C][=C][Ring1][=Branch1][Branch1][=Branch2] [C][=C][C][=C][C][=C][Ring1][=Branch1][C][=C][C][=C][C][=C][Ring1][=Branch1].[Br-1][END_SELFIES] ¡IDX_Solvents_MAP¿: Solvents: \$2\$: [START_SELFIES][C][C][O][C][C][END_SELFIES] ACTION_Sequence: ADD \$1\$ (2 liter, 57.2 g) ; ADD \$2\$ (500 ml) ; ADD \$2\$ ; STIR ; ADD \$3\$ (18.0 g) ; ADD \$4\$ (23.7 g) ; FILTER ; WASH with water ; WASH with sodium chloride ; WASH with water ; WASH with sodium chloride ; DRYSOLUTION over magnesium sulfate ; FILTER keep filtrate ; CONCENTRATE ; YIELD \$-1\$ (20.8 g).

### v3

@@ H.6 Quantum Mechanics Property Prediction
Instruction: What is the HOMO-LUMO gap of this molecule? The compound SELFIES sequence is: <SELFIES_compound>: [N][=C][O][C][=C][C][=Branch1][Ring2][=N][Ring1][=Branch1][C][#N] Property: 0.1487
H.7 Catalyst Prediction
The catalyst prediction task focuses on identifying the appropriate catalysts required to facilitate a given chemical reaction. Catalysts play a crucial role in modifying reaction pathways, lowering activation energy, and improving reaction efficiency without being consumed in the process. The input follows the SELFIES representation, a robust molecular encoding system designed for computational applications. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using “>>" (e.g., “reactant1.reactant2>>product"). This structured representation allows the model to process the reaction as a whole and infer the most suitable catalyst that enables the transformation. One of the primary challenges in catalyst prediction is understanding the nuanced role that catalysts play in different reaction mechanisms. Unlike reagents, which directly participate in the reaction, catalysts provide alternative pathways to enhance reaction kinetics. The prompt template is as follows.
@@ H.7 Catalyst Prediction
Instruction: Given this chemical reaction, what are some catalysts that could have been used? The reaction is <SELFIES_reactants>: [C][C][=C][C][=C][Branch1][Ring1][C][#N][C][=C][Ring1][Branch2][C][Branch1][C][F][Branch1][C][F][F].[O][=C][C][C][C][=Branch1][C][=O][N][Ring1][=Branch1][Br] <SELFIES_product>: [N][#C][C][=C][C][=C][Branch1][Ring1][C][Br][C][Branch1][=Branch2][C][Branch1][C][F][Branch1][C][F][F][=C][Ring1][N] <SELFIES_catalysts>:[O][=C][Branch1][#C][O][O][C][=Branch1][C][=O][C][=C][C][=C][C][=C][Ring1][=Branch1][C]$[=C][C]$ [=C][Ring1][=Branch1]
H.8 Solvent Prediction
The solvent prediction task focuses on identifying the appropriate solvents required for a given chemical reaction. Solvents play a crucial role in determining reaction efficiency, influencing factors such as solubility, reaction kinetics, selectivity, and stability of intermediates. To ensure a structured and machine-readable representation, the input follows the SELFIES format, a robust molecular encoding system designed for computational applications. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using the reaction separator “>>" (e.g., “reactant1.reactant2>>product"). This structured format allows the model to interpret the reaction context and infer the most suitable solvents required to facilitate the transformation. One of the key challenges in solvent prediction is understanding the diverse roles solvents play in different reaction mechanisms. The prompt template is as follows.
@@ H.8 Solvent Prediction
Instruction: Please propose potential solvents that might have been utilized in the provided chemical reaction. The reaction is <SELFIES_reactants>: [N][#C][C][=C][C][=C][Branch1][C][F][C][=C][Ring1][#Branch1].[O][C][C][C][N][C][Ring1][Branch1] <SELFIES_product>: [N][#C][C][=C][C][=C][Branch1][O][N][C][C][C][Branch1][C][O][C][Ring1][=Branch1][C][=C][Ring1][N] <SELFIES_solvents>: [O]
H.9 Yield Regression
The yield regression task focuses on estimating the proportion of the actual product obtained in a chemical reaction relative to its theoretical maximum. Reaction yield is a critical metric in organic synthesis, pharmaceutical manufacturing, and industrial chemistry, as it directly influences process efficiency, resource utilization, and cost-effectiveness. The input follows the SELFIES format, a robust molecular encoding system tailored for computational chemistry. The reaction is expressed as an equation where the SELFIES representations of the reactants are concatenated and separated from the product using the reaction separator “>>" (e.g., “reactant1.reactant2>>product"). This structured format provides a standardized input for the model, allowing it to interpret the reaction context and estimate the expected yield. One of the key challenges in yield prediction is capturing the complex interplay between reaction conditions, molecular stability, steric effects, and solvent or catalyst influences. The prompt template is as follows.
@@ H.9 Yield Regression
Instruction: Please propose potential solvents that might have been utilized in the provided chemical reaction. The reaction is <SELFIES_reactants>: [F][C][Branch1][C][F][Branch1][C][F][C][=C][C][=C][Branch1][C][Cl][C][=C][Ring1][#Branch1].[C][C][=C][C][=C][Branch1][C][N][C][=C][Ring1][#Branch1].[O][=S][=Branch1][C][=O][Branch2][Ring1][=Branch1][O][Pd][N][C][=C][C][=C][C][=C][Ring1][=Branch1][C][=C][C][=C][C][=C][Ring1][=Branch1][Ring1][=C][C][Branch1][C][F][Branch1][C][F][F].[C][O][C][=C][C][=C][Branch1][Ring1][O][C][C][Branch2][Ring2][=N][P][Branch2][Ring1][Branch1][C][C][C][C][C][Branch1][O][C][C][Branch1][Ring2][C][Ring1][=Branch1][C][Ring1][=Branch2][C][Ring1][#Branch2][C][C][C][C][C][Branch1][O][C][C][Branch1][Ring2][C][Ring1][=Branch1][C][Ring1][=Branch2][C][Ring1][#Branch2][=C][Ring2][Ring1][=N][C][=C][Branch1][=Branch1][C][Branch1][C][C][C][C][=C][Branch1][=Branch1][C][Branch1][C][C][C][C][=C][Ring1][N][C][Branch1][C][C][C].[C][N][C][C][C][N][C][C][C][N][=C][Ring1][#Branch2][Ring1][=Branch1].[C][C][O][C][=Branch1][C][=O][C][C][=C][Branch1][C][C][O][N][=Ring1][=Branch1] <SELFIES_product>: [C][C][=C][C][=C][Branch2][Ring1][Ring2][N][C][=C][C][=C][Branch1][=Branch2][C][Branch1][C][F][Branch1][C][F][F][C][=C][Ring1][#Branch2][C][=C][Ring1][P] Property: 0.1449
H.10 LogP Prediction
@@
Instruction: I am interested in the LogP of this molecule, could you tell me what it is? If uncertain, provide an estimate. Respond with the numerical value only. The molecule SELFIES sequence is: <SELFIES_compound>: [C][O][C][=C][C][=Branch1][=Branch2][=C][C][=Branch1][Ring2][=C][Ring1][=Branch1][O][C@H1][C@H1][C@@H1][Branch1][Branch1][C][O][Ring1][Branch1][C@@H1][Branch1][=Branch1][O][C][Ring1][#Branch1][=O][C][=C][C][=Branch1][=C][=C][Branch1][=Branch2][C][=Branch1][Ring2][=C][Ring1][=Branch1][O]$[C][O]$ Property: The LogP for the input molecule is 2.00.
H.11 Molecular Weight Prediction
@@
Instruction: Please provide me with the Molecular Weight value of this molecule. Determine the Molecular Weight value of this molecule. If uncertain, provide an estimate. Respond with the numerical value only. The molecule SELFIES sequence is: <SELFIES_compound>: [C][O][C][=C][Branch1][#Branch1][C][=C][C][=N][Ring1][=Branch1][C][=Branch1][C][=O][N][C][C][C][C][C][N][Branch1] [Branch1][C][C][Ring1][=Branch1][S][=Branch1][C][=O][=Branch1][C][=O][N][C][=Branch1][C][=O][N][C][C][C][C][C][C][Ring1][Branch1][C][=C][Ring1][Branch1] Property: The Molecular Weight for the input molecule is 491.60 g/mol.
H.12 Topological Polar Surface Area Prediction
@@
Instruction: I would like to know the Topological Polar Surface Area of this molecule, can you provide it? If uncertain, provide an estimate. Respond with the numerical value only. The compound SELFIES sequence is: <SELFIES_compound>:[C][C][=Branch2][=Branch1][=Branch2][=C][C][O][C][C][Branch1][O][C][Branch1][Ring2][O][Ring1][=Branch1][Branch1][C][C][C][O][C][C][C][C][Branch2][Ring2][#C][C][Branch2][Ring2][#Branch2][C][Branch2][Ring1][=Branch1][C][C][Branch1][P][C][Ring1][=Branch1][Branch1][O][C][Ring1][#Branch2][C][Ring2][Ring1][C][O][Ring1][Ring1][O][O][C][C][=C][Ring1][=N][N][C][=C][C][=C][C][=C][Ring1][=Branch2][Ring1][=Branch1][C]$[C]$ Property: The Topological Polar Surface Area for the input molecule is 96.50 Ų.
H.13 Description Q&A
@@
Instruction: What is the main component of Lobaplatin that gives it its anticancer properties? The compound SELFIES sequence is: <SELFIES_compound>: [C][C@@H1][Branch1][=Branch1][C][=Branch1][C][=O][O-1][O-1].[C][C][C@H1][Branch1][=Branch1][C@@H1][Ring1][Ring2][C][N][C][N].[Pt+2] Description: The main component of Lobaplatin that gives it its anticancer properties is the platinum cation (Pt+2).
H.14 Experimental Procedure Prediction
The experimental procedure prediction task focuses on generating a detailed, step-by-step protocol for conducting chemical experiments based on a given set of experimental components. This task plays a crucial role in automated synthesis planning, high-throughput experimentation, and robotic chemistry, where structured experimental procedures are essential for reproducibility, efficiency, and accuracy. The input is a mapping between experimental components and their corresponding SELFIES representations (e.g., “Reactants: \$index\$: SELFIES" …). The model is tasked with producing a structured sequence of operations that associates each component with the detailed steps of the experiment. The objective is to automate the chemical synthesis process by providing executable, structured experimental procedures. The prompt template is as follows.
@@ H.14 Experimental Procedure Prediction
Instruction: From the provided starting materials and target compound, generate the step-by-step experimental protocol. The Action Sequence: <IDX_Reactants_MAP>:Reactants:\$4\$:[START_SELFIES][O][=C][C][=C][C][=C][Branch1][O][O][C][C][C][O][C][C][O][Ring1][Branch1][C][=C][Ring1][=C][END_SELFIES] \$3\$: [START_SELFIES][C][C][Branch1][C][C][Branch1][C][C][O-1].[K+1][END_SELFIES] <IDX_Product_MAP>:Product:\$-1\$:[START_SELFIES][C][=C][C][=C][C][=C][Branch1][O][O][C][C][C][O][C][C][O][Ring1][Branch1][C][=C][Ring1][=C][END_SELFIES] <IDX_Catalysts_MAP>:Catalysts:\$1\$:[START_SELFIES][C][P+1][Branch1][=Branch2][C][=C][C][=C][C][=C][Ring1][=Branch1][Branch1][=Branch2] [C][=C][C][=C][C][=C][Ring1][=Branch1][C][=C][C][=C][C][=C][Ring1][=Branch1].[Br-1][END_SELFIES] <IDX_Solvents_MAP>: Solvents: \$2\$: [START_SELFIES][C][C][O][C][C][END_SELFIES] ACTION_Sequence: ADD \$1\$ (2 liter, 57.2 g) ; ADD \$2\$ (500 ml) ; ADD \$2\$ ; STIR ; ADD \$3\$ (18.0 g) ; ADD \$4\$ (23.7 g) ; FILTER ; WASH with water ; WASH with sodium chloride ; WASH with water ; WASH with sodium chloride ; DRYSOLUTION over magnesium sulfate ; FILTER keep filtrate ; CONCENTRATE ; YIELD \$-1\$ (20.8 g).
H.15 IUPAC2SELFIES
In the IUPAC2SELFIES task, the model is asked to generate the SELFIES representation of a molecule based on its IUPAC name. The IUPAC (International Union of Pure and Applied Chemistry) name is a standardized and systematic way to uniquely describe a molecule’s structure using rules for parent chains, functional groups, and substituents. To succeed in this task, the model must accurately understand these rules and produce the correct SELFIES representation accordingly.
  Template
[…]
Instruction: What is the molecular structure corresponding to this IUPAC name? The IUPAC name is: N-(1-amino-2-methylpropan-2-yl)-N-propylmorpholine-4-carboxamide <SELFIES_compound>: [C][C][C][N][Branch1][=N][C][=Branch1][C][=O][N][C][C][O][C][C][Ring1][=Branch1][C][Branch1][C][C][Branch1][C][C][C][N]
H.16 Text Guided Molecule Generation
In this task, the model is asked to generate a molecule in SELFIES notation based on a textual description. The description includes information such as functional group substitutions, structural features, and desired properties. The model must comprehend this complex information and design a molecule that satisfies the given requirements.
  Template
[…]
Instruction: Design a molecule according to the description provided. The description is: The molecule is a nitrile that is propionitrile in which one of the methyl hydrogens has been replaced by a phenyl group. It is a nitrile and a member of benzenes. It is functionally related to an acetonitrile. The molecule is a natural product found in Brassica napus, Brassica oleracea, and other organisms with data available. <SELFIES_compound>: [C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][C][C][C][#N]
H.17 Molecule Editing
The Molecule Editing task is concerned with optimizing one or more target properties of a given molecule, particularly the DRD2 score and QED. Specifically, the objective is to apply structural edits to a molecule with low DRD2 score and low QED in order to generate a variant exhibiting higher DRD2 score and QED. The model is tasked with optimizing the QED or DRD2 score of an input SELFIES molecule while maintaining some extent of structural similarity. It outputs the SELFIES of the optimized molecule. The objective is to optimize specific properties of a given molecule so as to yield novel biological insights. The prompt template is as follows.
  Template
[…]
Instruction: Develop a new molecular structure that enhances the QED score and keeps structural similarity to the initial molecule. The compound SELFIES sequence is: <SELFIES_compound_Input>: [C][C][=Branch1][C][=O][N][C][C][C][N][Branch2][Ring1][C][C][=C][C][=C][Branch1][Branch2][C][=Branch1][C][=O][C][C][#N][C][=C][Ring1][O][C][=Branch1][C][=O][O][Ring1][P] <SELFIES_compound>: [C][C][=Branch1][C][=O][N][C][C][C][N][Branch2][Ring1][Ring1][C][=C][C][=C][Branch1][=Branch2][C][Branch1][C][C][Branch1][C][C][C][C][=C][Ring1][#Branch2][C][=Branch1][C][=O][O][Ring1][S]
Appendix I Discussion on Generalist and Specialist
Here, we define the notions of generalist and specialist models. A generalist model refers to a single model with shared parameters $\theta$ that is trained across an entire set of tasks $\mathcal{T}=\{T_{i}\}$. In contrast, a specialist model assigns a distinct set of parameters $\theta_{i}$ to each individual task $T_{i}$.
Although the authors of InstructMol and HIGHT claim their models to be generalist in their respective papers, they in fact employ distinct LoRA adapters for different tasks, following the formulation $\theta_{i}=\theta_{0}+\theta_{\text{lora}_{i}}$, where $\theta_{0}$ denotes the pretrained LLM parameters and $\theta_{\text{lora}_{i}}$ represents the task-specific LoRA parameters for task $T_{i}$. As a result, these models do not satisfy the criteria for being true generalist models. In contrast, both PRESTO and Omni-Mol utilize a single parameter set $\theta$ shared across all tasks.
Appendix J Discussion on Continual Learning
Configuration Exact BLEU Levenshtein RDK MACCS Morgan Validity
Continual Learning Performance on Retrosynthesis
Retrosynthesis only 0.55 0.95 9.93 0.83 0.89 0.80 1.00
Forward + Retrosynthesis 0.50 0.94 11.01 0.77 0.82 0.76 1.00
Table 13: Continual Learning performance on retrosynthesis task before and after learning on forward prediction.
Instead of unified instruction tuning, which trains the model to learn all tasks simultaneously in a single procedure, it is also plausible to train the model sequentially on various tasks.
We conduct an experiment in a continual learning fashion: we first train the model on the retrosynthesis task and then use the trained model to learn forward prediction. We record the performance on the retrosynthesis task both before and after training on forward prediction. As shown in Table 13, after training on forward prediction, we observed a significant performance drop. This suggests that the model may forget previously learned knowledge and cannot benefit from earlier tasks. Therefore, in Omni-Mol, we train all tasks jointly.